Lynch Company manufactures and sells a single product. The following costs were incurred during the companys first year of oComplete this question by entering your answers in the tabs below. Req 1A Reg 1B Reg 2A Req 2B Compute the unit product cost.Prepare an income statement for the year. Assume that the company uses absorption costing. Lynch Company Absorption Costing IReq 1A Reg 1B Req 2A Req 2B Compute the unit product cost. Assume that the company uses variable costing Unit product cost $Req 1A Reg 1B Reg 2A Reg 2B Prepare an income statement for the year. Assume that the company uses variable costing Lynch Com
Answer #1
1a
Direct materials 6
Direct labor 9
Variable manufacturing overhead 3
Fixed manufacturing overhead 12 =300000/25000
Unit product cost 30
b
Absorption Costing Income Statement
Sales 1000000 =20000*50
Cost of goods sold 600000 =20000*30
Gross margin 400000
Selling and administrative expenses 270000 =190000+(20000*4)
Net operating income 130000
2a
Direct materials 6
Direct labor 9
Variable manufacturing overhead 3
Unit product cost 18
b
Variable Costing Income Statement
Sales 1000000 =20000*50
Variable expenses:
Variable Cost of goods sold 360000 =20000*18
Variable Selling and administrative expenses 80000 =20000*4
440000
Contribution margin 560000
Fixed expenses:
Fixed manufacturing overhead 300000
Fixed Selling and administrative expenses 190000
490000
Net operating income 70000